OpenStudy (anonymous):

A and B have coordinates –2 and 10 on a number line. If it is marked off into 10 congruent segments, which one of the following coordinates is NOT one of the marked points?

OpenStudy (anonymous):

please help me somebody i try to pass this class

OpenStudy (anonymous):

you didn't give any answer choices.

OpenStudy (anonymous):

a. –0.8 c. 1.6 b. –0.6 d. 2.8

OpenStudy (anonymous):

the distance between the points is 12 so dividing it into 10 segments gives each segment a length of 10/12 = 5/6 = 0.83. So any multiple of 5/6 will be a marked point. just pick which choice is not a multiple of 5/6.